# Nonlocal critical problems with mixed boundary conditions and nearly resonant perturbations

### v1

@@ Abstract.
where $(-\Delta)^{s}$, $s\in(1/2,1)$, is the spectral fractional Laplacian operator, $\Omega\subset\mathbb{R}^{N}$, $N>2s$, is a smooth bounded domain, $2_{s}^{*}=\frac{2N}{N-2s}$ denotes the critical fractional Sobolev exponent, $\lambda>0$ is a real parameter, $\nu$ is the outwards normal to $\partial\Omega$, $\Sigma_{\mathcal{D}}$, $\Sigma_{\mathcal{N}}$ are smooth (N1)(N-1)–dimensional submanifolds of $\partial\Omega$ such that $\Sigma_{\mathcal{D}}\cup\Sigma_{\mathcal{N}}=\partial\Omega$, $\Sigma_{\mathcal{D}}\cap\Sigma_{\mathcal{N}}=\emptyset$ and $\Sigma_{\mathcal{D}}\cap\overline{\Sigma}_{\mathcal{N}}=\Gamma$ is a smooth (N2)(N-2)–dimensional submanifold of $\partial\Omega$. By employing both a $\nabla$-theorem combined with a linking-type theorem, we prove the existence of multiple solutions when the parameter $\lambda$ is in a left neighborhood of a given eigenvalue of $(-\Delta)^{s}$.
@@ 1. Introduction
Assume $(\Omega_{1})-(\Omega_{4})$ and let $\Lambda_{k}$, $k\geq 2$, be an eigenvalue of the problem
@@
for every $\lambda\in\left(\Lambda_{k}-{\delta_{k}},\Lambda_{k}\right)$, ($P_{\lambda}$) has at least three solutions.
@@
Although $\nabla$-theorems have been previously applied to obtain multiplicity of solutions (cf. [1, 13, 14, 15, 18, 19]), this is, up to our knowledge, the first work where this tool is used to deal with critical problems. In this case, in addition to the usual geometric and compactness analysis, a careful analysis is required to obtain explicit estimates of some functional quantities related to ($P_{\lambda}$). To be more precise, the key points of our approach are the following. After performing a spectral decomposition of the underlying energy space, we first determine the precise energy threshold $\varepsilon_{\lambda}$ below which one obtains only trivial critical points for the energy functional $I_{\lambda}$ associated with ($P_{\lambda}$) restricted to a certain subspace (see Lemma 8 below). Regarding this step, the harmonic mean of eigenvalues adjacent to the eigenvalue $\Lambda_{k}$ plays a crucial role (see Remark 9), guaranteeing that $\varepsilon_{\lambda}$ is, in turn, lower than the compactness threshold $c^{*}$ of Palais-Smale sequences (being $c^{*}$ determined in Proposition 7). Next, we prove the compactness of sequences whose projections on the eigenspace associated with the eigenvalue $\Lambda_{k}$ and on its orthogonal satisfy certain conditions (see Lemma 10). This step is crucial to prove the validity of the so-called $\nabla$-condition, a major ingredient of the $\nabla$-theorem (see Definition 3). Lastly, we obtain a precise bound for $I_{\lambda}$ on the eigenspace spanned by the first k+m1k+m-1 eigenvalues, $m$ being the multiplicity of $\Lambda_{k}$ (see Lemma 12). The combination of these results allows us to apply the aforementioned $\nabla$-theorem (see Theorem 4) to deduce the existence of at least two solutions to ($P_{\lambda}$) for $\lambda$ close to every $\Lambda_{k}$, $k\geq 2$. We conclude by proving that, under our assumptions, $I_{\lambda}$ has the linking structure needed to apply the classical linking result by Rabinowitz (cf. [23]) and obtaining then a third solution, again for $\lambda$ near the eigenvalues of $(-\Delta)^{s}$. It is worth mentioning that the range of $\lambda$ in Theorem 1 follows the lines of that obtained in [9] by means of an abstract critical point theorem due to Bartolo, Benci and Fortunato (cf. [3]).
@@ 3. Linking geometry, $\nabla$- and $(PS)$-conditions
Now we address the fulfillment of the $\nabla$-condition for $I_{\lambda}$. This will require some preliminary steps. First, we prove the $(PS)_{c}$ for suitable levels $c\in\mathbb{R}$.
@@ 4. Proof of Theorem 1
for every $\lambda\in(\Lambda_{k}-\delta_{k},\Lambda_{k})$.
To conclude the proof, we show that, if $\lambda\in(\Lambda_{k-1},\Lambda_{k})$, $I_{\lambda}$ has an additional linking structure which allows us to deduce the existence of a third solution. Indeed, splitting $H_{\Sigma_{\mathcal{D}}}^{s}(\Omega)=\mathbb{H}_{k-1}\oplus\mathbb{P}_{k-1}$ and retracing the proof of Proposition 6, we deduce the existence of $\varrho,\beta>0$ such that has
$I_{\lambda}(u)\geq\beta,\qquad\text{for all }u\in\partial B_{\varrho}(\mathbb{P}_{k-1}).$
On the other hand, by (3.2), we have ${I_{\lambda}}_{|\mathbb{H}_{k-1}}\leq 0$. By (3.3) one has also $I_{\lambda}(u)\leq 0$ for $u\in\mathbb{H}_{k-1}\oplus\text{span}\{\varphi_{k}\}=\mathbb{H}_{k}$ and $\left\|u\right\|_{H_{\Sigma_{\mathcal{D}}}^{s}}\geq\bar{R}$, for some $\bar{R}>\varrho$. Then, for a sufficiently large $R>0$, setting
$Q:=(B_{R}(H_{\Sigma_{\mathcal{D}}}^{s}(\Omega))\cap\mathbb{H}_{k-1})\oplus\{r\varphi_{k},\ r\in(0,R)\},$
one has ${I_{\lambda}}_{|\partial Q}\leq 0$, where $\partial Q$ is the boundary of $Q$ relative to $\mathbb{H}_{k-1}\oplus\ \text{span}\{\varphi_{k}\}$). So, by [23, Theorem 5.3] (see also [23, Remark 5.5]), $I_{\lambda}$ has a third critical point $u_{3}$ satisfying
$I_{\lambda}(u_{3})\geq\inf\limits_{u\in\partial B_{\varrho}(\mathbb{P}_{k-1})}I_{\lambda}(u)\geq\beta.$
Moreover, $u_{3}\neq u_{i}$, $i=1,2$, because, on account of 4.6 and Lemma 12, for $\lambda$ close to $\Lambda_{k}$ one has
$I_{\lambda}(u_{i})\leq\sup\limits_{u\in\mathbb{H}_{k+m-1}}I_{\lambda}(u)<\inf\limits_{u\in\partial B_{\varrho}(\mathbb{P}_{k-1})}I_{\lambda}(u).$
Then, Theorem 1 is proved. ∎

### v2

@@ Abstract.
where $(-\Delta)^{s}$, $s\in(1/2,1)$, is the spectral fractional Laplacian operator, $\Omega\subset\mathbb{R}^{N}$, $N>2s$, is a smooth bounded domain, $2_{s}^{*}=\frac{2N}{N-2s}$ denotes the critical fractional Sobolev exponent, $\lambda>0$ is a real parameter, $\nu$ is the outwards normal to $\partial\Omega$, $\Sigma_{\mathcal{D}}$, $\Sigma_{\mathcal{N}}$ are smooth (N1)(N-1)–dimensional submanifolds of $\partial\Omega$ such that $\Sigma_{\mathcal{D}}\cup\Sigma_{\mathcal{N}}=\partial\Omega$, $\Sigma_{\mathcal{D}}\cap\Sigma_{\mathcal{N}}=\emptyset$ and $\Sigma_{\mathcal{D}}\cap\overline{\Sigma}_{\mathcal{N}}=\Gamma$ is a smooth (N2)(N-2)–dimensional submanifold of $\partial\Omega$. By employing a $\nabla$-theorem we prove the existence of multiple solutions when the parameter $\lambda$ is in a left neighborhood of a given eigenvalue of $(-\Delta)^{s}$.
@@ 1. Introduction
Assume $(\Omega_{1})-(\Omega_{4})$ and let $\Lambda_{k}$, $k\geq 2$, be an eigenvalue of multiplicity $m\in\mathbb{N}$ of the problem
@@
for every $\lambda\in\left(\Lambda_{k}-{\delta_{k}},\Lambda_{k}\right)$, ($P_{\lambda}$) has at least two solutions.
@@
Although $\nabla$-theorems have been previously applied to obtain multiplicity of solutions (cf. [1, 13, 14, 15, 18, 19]), this is, up to our knowledge, the first work where this tool is used to deal with critical problems. In this case, in addition to the usual geometric and compactness analysis, a careful analysis is required to obtain explicit estimates of some functional quantities related to ($P_{\lambda}$). To be more precise, the key points of our approach are the following. After performing a spectral decomposition of the underlying energy space, we first determine the precise energy threshold $\varepsilon_{\lambda}$ below which one obtains only trivial critical points for the energy functional $I_{\lambda}$ associated with ($P_{\lambda}$) restricted to a certain subspace (see Lemma 8 below). Regarding this step, the harmonic mean of eigenvalues adjacent to the eigenvalue $\Lambda_{k}$ plays a crucial role (see Remark 9), guaranteeing that $\varepsilon_{\lambda}$ is, in turn, lower than the compactness threshold $c^{*}$ of Palais-Smale sequences (being $c^{*}$ determined in Proposition 7). Next, we prove the compactness of sequences whose projections on the eigenspace associated with the eigenvalue $\Lambda_{k}$ and on its orthogonal satisfy certain conditions (see Lemma 10). This step is crucial to prove the validity of the so-called $\nabla$-condition, a major ingredient of the $\nabla$-theorem (see Definition 3). Lastly, we obtain a precise bound for $I_{\lambda}$ on the eigenspace spanned by the first k+m1k+m-1 eigenvalues, $m$ being the multiplicity of $\Lambda_{k}$ (see Lemma 12). The combination of these results allows us to apply the aforementioned $\nabla$-theorem (see Theorem 4) to deduce the existence of at least two solutions to ($P_{\lambda}$) for $\lambda$ close to every $\Lambda_{k}$, $k\geq 2$. It is worth mentioning that the range of $\lambda$ in Theorem 1 follows the lines of that obtained in [9] by means of an abstract critical point theorem due to Bartolo, Benci and Fortunato (cf. [3]).
@@ 3. Linking geometry, $\nabla$- and $(PS)$-conditions
Now we address the fulfillment of the $\nabla$-condition for $I_{\lambda}$. This will require some preliminary steps. First, we prove the $(PS)_{c}$ condition for suitable levels $c\in\mathbb{R}$.
@@ 4. Proof of Theorem 1
for every $\lambda\in(\Lambda_{k}-\delta_{k},\Lambda_{k})$. Then, Theorem 1 is proved. ∎